# Dual-Prototype Disentanglement: A Context-Aware Enhancement Framework for Time Series Forecasting

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
  • We introduce a novel context-aware enhancement framework based on dual-prototype disentanglement. Unlike existing methods, our framework equips models with the capability for context-aware and pattern disentanglement, thereby directly mitigating the inherent limitation of static representations and enhancing their ability to dynamically modeling intricate temporal patterns.
@@ 2.2 Enhancement Strategies for Time Series Forecasting
Beyond architectural innovations, several efficient and model-agnostic strategies have been proposed to enhance predictive performance. Some methods employ decomposition strategies as a strong inductive bias to guide model in handling diverse temporal patterns more effectively (Huang et al., 2025; Wang et al., 2024a). Autoformer (Wu et al., 2021) decomposes time series into seasonal and trend components, allowing other models to handle input from a clearer perspective. TimesNet (Wu et al., 2023a) learns frequency-adaptive periods in the frequency domain to explicitly model intra- and inter-period variations. Another strategy focuses on model-agnostic methods (Liu et al., 2025c; Ning et al., 2025). RAFT (Han et al., 2025) employs retrieval augmented generation to retrieves from historical samples, providing the model with relevant information. HCAN (Sun et al., 2024) assigns class labels for timesteps to equip models with contextual-aware capability. Furthermore, specialized loss function designs have been designed to improve robustness against distribution shifts and intricate patterns (Wang et al., 2025a; Liu et al., 2025b). DBLoss (Qiu et al., 2025b) decomposes both predictions and ground truth to calculate the loss for each components separately. PSLoss (Kudrat et al., 2025) enhances structural alignment by comparing time series at the patch level with statistical properties. While these strategies improve forecasting performance, they still fail to dynamically disentangle complex, intertwined patterns in a context-aware manner. Different from them, our proposed DPAD framework introduces a learnable dual-prototype bank and adaptively fuses them through a context-aware routing mechanism, thereby enabling dynamic and interpretable enhancement of the backbone model’s representations with high efficiency.
@@ 3.2 Dynamic Dual-Prototype Bank
The Dynamic Dual-Prototype bank (DDP) serves as the cornerstone of our framework, designed to explicitly disentangle and memorize two distinct categories of temporal patterns: prevalent common patterns and critical yet infrequent rare patterns. Formally, the DDP consists of two separate prototype sets: a Common Pattern Bank $\mathcal{B}_{c}$, which comprise sequences synthesized from a combination of fundamental trends and seasonal patterns, and a Rare Pattern Bank $\mathcal{B}_{r}$, dedicated to memorizing long-tailed, sudden events and some other temporal noise.
[…]
@@
In contrast, $\mathbf{s}_{r}^{j}$ is initialized from Gaussian distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I})$, where $\sigma^{2}$ is a small variance controlling the magnitude of initialization noise. This minimal random initialization ensures that rare prototypes begin from an nearly unbiased state, allowing them to remain flexible and exclusively adapt to capture rare patterns during training, without being predisposed to common modes. Notably, all of sequence sets would be transformed into learnable parameters.
@@ 3.3 Dual-Path Context-Aware Routing
$\mathbf{z}_{c}=\sum_{k\in\mathcal{I}_{c}}\omega_{c}^{k}\cdot\mathcal{B}_{c}[k],\quad\mathbf{z}_{r}=\omega_{r}\cdot\mathcal{B}_{r}[\mathcal{I}_{r}],$ (10)
@@ 3.4 Disentanglement-Guided Loss
Then we propose a Rarity Preservation Loss $\mathcal{L}_{\text{rare}}$. With the contrastive restriction, $\mathcal{L}_{\text{rare}}$ prevent the prototypes in the rare bank from being contaminated or forgotten due to common patterns, thereby ensuring their persistent memory for rare patterns:
@@ 4.1 Experimental Setup
Baselines. To verify the performance and effectiveness of our DPAD, we select five state-of-art (SOTA) time series forecasting models with diverse architectures, including transformer-based models: iTransformer (Liu et al., 2024a), TimeXer (Wang et al., 2024c), TimeBridge (Liu et al., 2025a); MLP-based models: DLinear (Zeng et al., 2023); and CNN-based models: TimesNet (Wu et al., 2023a).
@@
Table 1: Forecasting results for long-term and short-term. The look-back length is fixed to 96 for all baselines. All the results are averaged from 4 different prediction lengths, that is $\{12,24,48,96\}$ for PEMS and $\{96,192,336,720\}$ for others. The better results are highlighted in bold. See Section E.1 for full results.
@@ 4.2 Main Results
We present the main results for both long-term and short-term forecasting on seven real-world datasets for five state-of-art forecasting models in Table 1, where the lower MSE and MAE indicate better forecasting performance. Our proposed DPAD framework demonstrates consistent improvements across all backbone models, achieving the lowest MSE and MAE in most cases. Notably, DPAD achieves an average MSE reduction of $12.6\%$ for DLinear and $9.3\%$ for iTransformer, and follows a similar trend for other backbones. The performance gains are particularly substantial on datasets with inherent complexities such as Traffic and PEMS. These reduction demonstrates the effectiveness of our model-agnostic design, which equips diverse architectures with context-aware and pattern disentanglement capability without modifying their core structure. The above observations highlight the efficacy of DPAD in mitigating the inherent limitations of static representation in forecasting models.
@@ 4.3 Ablation Studies
Table 2: Ablation study on DDP to validate the efficacy of Common Pattern Bank and Rare Pattern Bank. $\checkmark$ and $\times$ indicate with and without certain components respectively. All the results are averaged from 4 different prediction lengths $\{96,192,336,720\}$, see Section E.2 for full results.
@@ 4.4 Zero-Shot Forecasting
To further evaluate the generalization and robustness of our DPAD on unseen datasets, we conduct zero-shot forecasting experiments. Following by prior works(Kudrat et al., 2025; Jin et al., 2024), we sequentially use ETTh1, ETTh2, ETTm1, and ETTm2 as source datasets, and the remaining datasets as target datasets.
@@
Table 5 shows zero-shot results with four backbone models, which consistently demonstrate that models enhanced with our DPAD framework achieve superior performance in most cases. These improvements indicate that the disentangled pattern memory in DPAD provides a more transferable and robust temporal representation. Specifically, the common prototypes in DDP learn domain-invariant, prevalent patterns that serve as reusable components across different datasets. Meanwhile, the DPC mechanism enables the model to dynamically integrate these prototypes, and activate sparse rare prototypes when necessary, which could account for novel, target-specific variations without requiring retraining. Consequently, these zero-shot results underscore that DPAD not only enhances forecasting performance for in-distribution data, but learns temporal patterns with more generalization.
@@ 4.5 Model Analysis
Efficiency Analysis. We conduct the efficiency analysis for the backbone model with and without DPAD. As shown in Table 6, DPAD introduces minimal overhead for most backbone models. The relative increases in running time and memory footprint are generally modest, which is below 10% for most Transformer or CNN-based backbones. While a lightweight model like DLinear shows a higher relative increase due to its extremely small base cost, the absolute added overhead remains minimal. This demonstrates DPAD is a practical and efficient enhancement module, delivering significant performance improvements while incurring only a small fraction of additional computational cost.
@@ 5 Conclusion
In this paper, we address a critical limitation of static and averaged representations in deep time series forecasting, which hinders model adaptability to intricate and dynamic contexts. We propose the Dual-Prototype Adaptive Disentanglement (DPAD) framework, a model-agnostic auxiliary method that enhances backbone model through a Dynamic Dual-Prototype bank (DDP), Dual-Path Context-aware routing (DPC), and a Disentanglement-Guided Loss (DGLoss). Extensive experiments validate that DPAD consistently improves forecasting performance while introduces minimal computational overhead. Future work may explore more interpretable routing mechanisms and novel prototype paradigm with multi-modal representations.
@@ Appendix A Datasets
We conduct long-term forecasting experiments on 6 real-world datasets, including (1) ETT (ETTh1, ETTh2, ETTm1, ETTm2) contains 7 features of electricity transformer data from July 2016 to July 2018, which sampled at hourly (ETTh1, ETTh2) and 15-minute (ETTm1, ETTm2) intervals. (2) Electricity (ECL) records hourly electricity consumption data of 321 clients from 2012 to 2014. (3) Exchange collects the daily exchange-rate data from eight different countries. (4) Solar-Energy contains the solar power production of 137 PV plants in 2006, which is sampled every 10 minutes. (5) Weather includes 21 meteorological factors collected every 10 minutes from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. (6) Traffic contains hourly road occupancy rates from 862 sensors on San Francisco freeways. Meanwhile, we conduct short-term forecasting experiments on PEMS (PEMS03, PEMS04, PEMS07, PEMS08), which contains the public traffic network data in California with a 5-minute interval. The details of datasets are listed in Table 7.
@@ C.1 Size of Prototype Banks
The sizes of the Common Pattern Bank $\mathcal{B}_{c}$ and Rare Pattern Bank $\mathcal{B}_{r}$ determine the representational capacity and specialization degree of our Dual‑Prototype Bank. To analyze their impact, we set $\mathcal{B}_{c}\in\{32,64,128,256\}$ and $\mathcal{B}_{r}\in\{8,12,24,32\}$ respectively on Weather and ETTh2 dataset. As shown in Figure 5, the results remains stable across a broad range of sizes, indicating that DPAD is not highly sensitive to the size of prototype banks. Extremely small banks may limit the ability to capture pattern diversity, while very large banks offer no further gain and may slightly degrade performance due to increased optimization difficulty. The above observation confirm that DPAD is robust across various cases without requiring delicate tuning.
@@ C.2 Embedding Dimension
The embedding dimension determines the richness of feature representations in the unified projection space. We vary $d_{\text{model}}\in\{64,128,256,512,1024\}$ on Weather and ETTh1 dataset with four backbone models to evaluate its impact. As shown in Figure 6, in most cases DPAD exhibits low sensitivity to the choice of embedding dimension. Performance does not degrade significantly even at relatively small dimensions, while larger dimensions provide only marginal gains. So we choose the optimal dimension at 128/256 in most cases. The above observation indicates that our dual-prototype architecture and routing mechanism are effective at capturing and utilizing pattern information without requiring a high dimension latent space. This property further simplifies the deployment of DPAD as a plug-and-play module.
@@
Figure 6: Hyperparameter sensitivity analysis on embedding dimension. For each backbone, forecasting results is on the Weather (left) and ETTh1 (right) datasets. The look-back length length is fixed to 96.
@@ C.3 Weights of DGLoss
The weight $\lambda_{\text{sep}},\lambda_{\text{rare}},\lambda_{\text{div}}$ of the DGLoss control the strength of separation, rarity preservation, and diversity of temporal prototypes. To analyze their individual effects, we vary them in 0.0012.00.001-2.0 on electricity and weather datasets. As shown in Figure 7,
@@
Figure 7: Hyperparameter sensitivity analysis on weights of DGLoss. Forecasting results is on the ECL and Weather datasets, and we use TimeBridge as backbone model. The look-back length length is fixed to 96.
Appendix D Forecasting Visualization
@@
Appendix E Full Results
E.1 Main experiments
we evaluate DPAD on seven real-world TSF benchmarks spanning diverse domains. Table 8 shows the full results of long-term forecasting tasks on ETT (ETTh1, ETTh2 , ETTm1, ETTm2), Electricity, Exchange, Solar, Weather, and Traffic datasets. Table 9 contains the full results of short-term forecasting tasks on PEMS (PEMS03, PEMS04, PEMS05, PEMS06) datasets. From full results, we can see that DPAD achieves consistent improvements on most backbones and datasets.

### v2

@@ Abstract
Time Series Forecasting, Context-Aware, Pattern Disentanglement
@@ 1 Introduction
  • We introduce a novel context-aware enhancement framework based on dual-prototype disentanglement. Unlike existing methods, our framework equips models with the capability for context-aware and pattern disentanglement, thereby directly mitigating the inherent limitation of static representations and enhancing their ability to dynamically model intricate temporal patterns.
@@ 2.2 Enhancement Strategies for Time Series Forecasting
Beyond architectural innovations, several efficient and model-agnostic strategies have been proposed to enhance predictive performance. Some methods employ decomposition strategies as a strong inductive bias to guide model in handling diverse temporal patterns more effectively (Huang et al., 2025; Wang et al., 2024a). Autoformer (Wu et al., 2021) decomposes time series into seasonal and trend components, allowing other models to handle input from a clearer perspective. TimesNet (Wu et al., 2023a) learns frequency-adaptive periods in the frequency domain to explicitly model intra- and inter-period variations. Another strategy focuses on model-agnostic methods (Liu et al., 2025c; Ning et al., 2025). RAFT (Han et al., 2025) employs retrieval augmented generation to retrieve from historical samples, providing the model with relevant information. HCAN (Sun et al., 2024) assigns class labels for timesteps to equip models with contextual-aware capability. Furthermore, specialized loss function designs have been designed to improve robustness against distribution shifts and intricate patterns (Wang et al., 2025a; Liu et al., 2025b). DBLoss (Qiu et al., 2025b) decomposes both predictions and ground truth to calculate the loss for each component separately. PSLoss (Kudrat et al., 2025) enhances structural alignment by comparing time series at the patch level with statistical properties. While these strategies improve forecasting performance, they still fail to dynamically disentangle complex, intertwined patterns in a context-aware manner. Different from them, our proposed DPAD framework introduces a learnable dual-prototype bank and adaptively fuses them through a context-aware routing mechanism, thereby enabling dynamic and interpretable enhancement of the backbone model’s representations with high efficiency.
@@ 3.2 Dynamic Dual-Prototype Bank
The Dynamic Dual-Prototype bank (DDP) serves as the cornerstone of our framework, designed to explicitly disentangle and memorize two distinct categories of temporal patterns: prevalent common patterns and critical yet infrequent rare patterns. Formally, the DDP consists of two separate prototype sets: a Common Pattern Bank $\mathcal{B}_{c}$, which comprises sequences synthesized from a combination of fundamental trends and seasonal patterns, and a Rare Pattern Bank $\mathcal{B}_{r}$, dedicated to memorizing long-tailed, sudden events and some other temporal noise.
Specifically, we generate two groups of base sequences $\{\mathbf{s}_{c}^{i}\}$ and rare sequences $\{\mathbf{s}_{r}^{j}\}$ using distinct prior strategies to reflect their different pattern characteristics:
@@
In contrast, $\mathbf{s}_{r}^{j}$ is initialized from a Gaussian distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I})$, where $\sigma^{2}$ is a small variance controlling the magnitude of initialization noise. This minimal random initialization ensures that rare prototypes begin from a nearly unbiased state, allowing them to remain flexible and exclusively adapt to capture rare patterns during training, without being predisposed to common modes. Notably, all of sequence sets would be transformed into learnable parameters.
@@ 3.3 Dual-Path Context-Aware Routing
$\mathbf{z}_{c}=\sum_{k\in\mathcal{I}_{c}}\omega_{c}^{k}\cdot\mathcal{B}_{c}[\mathcal{I}_{c}],\quad\mathbf{z}_{r}=\omega_{r}\cdot\mathcal{B}_{r}[\mathcal{I}_{r}],$ (10)
@@ 3.4 Disentanglement-Guided Loss
Then we propose a Rarity Preservation Loss $\mathcal{L}_{\text{rare}}$. With the contrastive restriction, $\mathcal{L}_{\text{rare}}$ prevents the prototypes in the rare bank from being contaminated or forgotten due to common patterns, thereby ensuring their persistent memory for rare patterns:
@@ 4.1 Experimental Setup
Baselines. To verify the performance and effectiveness of our DPAD, we select five state-of-the-art (SOTA) time series forecasting models with diverse architectures, including transformer-based models: iTransformer (Liu et al., 2024a), TimeXer (Wang et al., 2024c), TimeBridge (Liu et al., 2025a); MLP-based models: DLinear (Zeng et al., 2023); and CNN-based models: TimesNet (Wu et al., 2023a).
@@
Table 1: Forecasting results for long-term and short-term. The look-back length is fixed to 96 for all baselines. All the results are averaged from 4 different prediction lengths, that is $\{12,24,48,96\}$ for PEMS and $\{96,192,336,720\}$ for others. The better results are highlighted in bold. See Section F.1 for full results.
@@ 4.2 Main Results
We present the main results for both long-term and short-term forecasting on seven real-world datasets for five state-of-art forecasting models in Table 1, where the lower MSE and MAE indicate better forecasting performance. Our proposed DPAD framework demonstrates consistent improvements across all backbone models, achieving the lowest MSE and MAE in most cases. Notably, DPAD achieves an average MSE reduction of $12.6\%$ for DLinear and $9.3\%$ for iTransformer, and follows a similar trend for other backbones. The performance gains are particularly substantial on datasets with inherent complexities such as Traffic and PEMS. This reduction demonstrates the effectiveness of our model-agnostic design, which equips diverse architectures with context-aware and pattern disentanglement capability without modifying their core structure. The above observations highlight the efficacy of DPAD in mitigating the inherent limitations of static representation in forecasting models.
@@ 4.3 Ablation Studies
Table 2: Ablation study on DDP to validate the efficacy of Common Pattern Bank and Rare Pattern Bank. $\checkmark$ and $\times$ indicate with and without certain components respectively. All the results are averaged from 4 different prediction lengths $\{96,192,336,720\}$, see Section F.2 for full results.
@@ 4.4 Zero-Shot Forecasting
To further evaluate the generalization and robustness of our DPAD on unseen datasets, we conduct zero-shot forecasting experiments. Following prior works(Kudrat et al., 2025; Jin et al., 2024), we sequentially use ETTh1, ETTh2, ETTm1, and ETTm2 as source datasets, and the remaining datasets as target datasets.
@@
Table 5 shows zero-shot results with four backbone models, which consistently demonstrate that models enhanced with our DPAD framework achieve superior performance in most cases. These improvements indicate that the disentangled pattern memory in DPAD provides a more transferable and robust temporal representation. Specifically, the common prototypes in DDP learn domain-invariant, prevalent patterns that serve as reusable components across different datasets. Meanwhile, the DPC mechanism enables the model to dynamically integrate these prototypes, and activate sparse rare prototypes when necessary, which could account for novel, target-specific variations without requiring retraining. Consequently, these zero-shot results underscore that DPAD not only enhances forecasting performance for in-distribution data, but also learns temporal patterns with more generalization.
@@ 4.5 Model Analysis
Efficiency Analysis. We conduct the efficiency analysis for the backbone model with and without DPAD. As shown in Table 6, DPAD introduces minimal overhead for most backbone models. The relative increases in running time and memory footprint are generally modest, which are below 10% for most Transformer or CNN-based backbones. While a lightweight model like DLinear shows a higher relative increase due to its extremely small base cost, the absolute added overhead remains minimal. This demonstrates DPAD is a practical and efficient enhancement module, delivering significant performance improvements while incurring only a small fraction of additional computational cost.
@@ 5 Conclusion
In this paper, we address a critical limitation of static and averaged representations in deep time series forecasting, which hinders model adaptability to intricate and dynamic contexts. We propose the Dual-Prototype Adaptive Disentanglement (DPAD) framework, a model-agnostic auxiliary method that enhances the backbone model through a Dynamic Dual-Prototype bank (DDP), Dual-Path Context-aware routing (DPC), and a Disentanglement-Guided Loss (DGLoss). Extensive experiments validate that DPAD consistently improves forecasting performance while introduces minimal computational overhead. Future work may explore more interpretable routing mechanisms and a novel prototype paradigm with multi-modal representations.
@@ Appendix A Datasets
We conduct long-term forecasting experiments on 6 real-world datasets, including (1) ETT (ETTh1, ETTh2, ETTm1, ETTm2) contains 7 features of electricity transformer data from July 2016 to July 2018, which was sampled at hourly (ETTh1, ETTh2) and 15-minute (ETTm1, ETTm2) intervals. (2) Electricity (ECL) records hourly electricity consumption data of 321 clients from 2012 to 2014. (3) Exchange collects the daily exchange-rate data from eight different countries. (4) Solar-Energy contains the solar power production of 137 PV plants in 2006, which is sampled every 10 minutes. (5) Weather includes 21 meteorological factors collected every 10 minutes from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. (6) Traffic contains hourly road occupancy rates from 862 sensors on San Francisco freeways. Meanwhile, we conduct short-term forecasting experiments on PEMS (PEMS03, PEMS04, PEMS07, PEMS08), which contains the public traffic network data in California with a 5-minute interval. The details of datasets are listed in Table 7.
@@ C.1 Size of Prototype Banks
The sizes of the Common Pattern Bank $\mathcal{B}_{c}$ and Rare Pattern Bank $\mathcal{B}_{r}$ determine the representational capacity and specialization degree of our Dual‑Prototype Bank. To analyze their impact, we set $\mathcal{B}_{c}\in\{32,64,128,256\}$ and $\mathcal{B}_{r}\in\{8,12,24,32\}$ respectively on the Weather and ETTh2 dataset. As shown in Figure 5, the results remain stable across a broad range of sizes, indicating that DPAD is not highly sensitive to the size of prototype banks. Extremely small banks may limit the ability to capture pattern diversity, while very large banks offer no further gain and may slightly degrade performance due to increased optimization difficulty. The above observation confirms that DPAD is robust across various cases without requiring delicate tuning.
@@ C.2 Embedding Dimension
The embedding dimension determines the richness of feature representations in the unified projection space. We vary $d_{\text{model}}\in\{64,128,256,512,1024\}$ on the Weather and ETTh1 dataset with four backbone models to evaluate its impact. As shown in Figure 6, in most cases DPAD exhibits low sensitivity to the choice of embedding dimension. Performance does not degrade significantly even at relatively small dimensions, while larger dimensions provide only marginal gains. So we choose the optimal dimension at 128/256 in most cases. The above observation indicates that our dual-prototype architecture and routing mechanism are effective at capturing and utilizing pattern information without requiring a high-dimension latent space. This property further simplifies the deployment of DPAD as a plug-and-play module.
@@
Figure 6: Hyperparameter sensitivity analysis on embedding dimension. For each backbone, forecasting results are on the Weather (left) and ETTh1 (right) datasets. The look-back length length is fixed to 96.
@@ C.3 Weights of DGLoss
The weight $\lambda_{\text{sep}},\lambda_{\text{rare}},\lambda_{\text{div}}$ of the DGLoss controls the strength of separation, rarity preservation, and diversity of temporal prototypes. To analyze their individual effects, we vary them in 0.0012.00.001-2.0 on electricity and weather datasets. As shown in Figure 7,
@@
Figure 7: Hyperparameter sensitivity analysis on weights of DGLoss. Forecasting results are on the ECL and Weather datasets, and we use TimeBridge as backbone model. The look-back length length is fixed to 96.
Appendix D More Results with other Enhancement Strategies
To further validate the effectiveness of DPAD, we compare it against other representative model-agnostic enhancement strategies. RAFT (Han et al., 2025) employs retrieval augmented generation to enhance performance from historical samples. HCAN(Sun et al., 2024) proposes a hierarchy-aware attention module supported by uncertainty-aware classifiers and a consistency loss to alleviate issues caused by the boundary effects during the classification of timesteps. FreDF (Wang et al., 2025a) bypasses the complexity of label autocorrelation by learning to forecast in the frequency domain. As shown in Table 8, DPAD consistently achieves superior performance in most cases across various datasets and forecasting horizons. This is due to its unified framework for context-aware and pattern disentanglement, which addresses the core limitations of static representation more directly than other strategies. Thus DPAD demonstrates consistently strong generalization across diverse forecasting scenarios.
Table 8: Forecasting results compared with other enhancement strategies on Electricity, ETTh2, and Weather datasets. We use iTransformer as backbone for HCAN, FreDF, and DPAD. The best results are highlighted in bold, and the second-best results are highlighted in underline.
Dataset Electricity ETTh2 Weather
Forecast length 96 192 336 720 96 192 336 720 96 192 336 720
ori MSE 0.148 0.167 0.179 0.208 0.301 0.380 0.423 0.431 0.176 0.225 0.281 0.361
(2024a) MAE 0.241 0.248 0.271 0.298 0.350 0.399 0.431 0.447 0.216 0.257 0.299 0.353
RAFT MSE 0.168 0.172 0.182 0.212 0.288 0.384 0.432 0.436 0.190 0.241 0.294 0.370
(2025) MAE 0.269 0.270 0.281 0.304 0.341 0.402 0.443 0.460 0.237 0.281 0.323 0.374
HCAN MSE 0.588 0.484 0.466 0.440 0.776 1.208 1.123 1.857 0.219 0.274 0.307 0.353
(2024) MAE 0.603 0.530 0.516 0.499 0.656 0.813 0.787 1.119 0.301 0.348 0.371 0.400
FreDF MSE 0.158 0.166 0.188 0.225 0.290 0.373 0.418 0.426 0.171 0.224 0.282 0.359
(2025a) MAE 0.248 0.259 0.277 0.308 0.342 0.394 0.429 0.443 0.209 0.255 0.298 0.350
+ DPAD MSE 0.147 0.162 0.172 0.201 0.293 0.374 0.416 0.422 0.170 0.223 0.279 0.355
MAE 0.234 0.251 0.271 0.289 0.345 0.393 0.425 0.435 0.208 0.256 0.295 0.347
Appendix E Forecasting Visualization
@@
Appendix F Full Results
F.1 Main Experiments
We evaluate DPAD on seven real-world TSF benchmarks spanning diverse domains. Table 9 shows the full results of long-term forecasting tasks on ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, Solar, Weather, and Traffic datasets. Table 10 contains the full results of short-term forecasting tasks on PEMS (PEMS03, PEMS04, PEMS05, PEMS06) datasets. From the full results, we can see that DPAD achieves consistent improvements on most backbones and datasets.